OpenStudy (anonymous):

Write the number 12 as a product of prime factors. ?

OpenStudy (anonymous):

2 X 3 X 2

OpenStudy (anonymous):

\[ 2^2 \times 3 \]

OpenStudy (anonymous):

the fancier method? Ah, I see.

OpenStudy (anonymous):

12 = 2 x 2 x 3

OpenStudy (anonymous):

not fancier that'w the way prime factorization is define I guess .. ;)

OpenStudy (anonymous):

How about 248?

OpenStudy (anonymous):

\[ 248 = 2^3 \times 31 \]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Glad to help.